lnx - ln(x + 2) = 3
This is what I tried:
lnx - ln(x+2) = 3
distributed
lnx - lnx - ln2 = 3
these cancel out leaving...
-LN2 = 3
DIVIDE BOTH SIDES BY -LN2
the answer seems wrong.

Answer by nerdybill(7384)

Review log rules:
.
lnx - ln(x + 2) = 3
ln [(x)/(x+2)] = 3
(x)/(x+2) = e^3
x = e^3(x+2)
x = e^3x+2e^3
x-e^3x = 2e^3
x(1-e^3) = 2e^3
x = 2e^3/(1-e^3)